Find (a) a basis for and (b) the dimension of the solution space of the 

homogeneous system of linear equations.

1.

x−2y +3z= 0

−3𝑥x +6𝑦y -9z = 0


1
Expert's answer
2021-06-14T12:19:57-0400

Let us write the system of equations in matrix form "A \bold x = 0", where "A = \begin{pmatrix}
1 & -2 & 3\\
-3 & 6 & -9
\end{pmatrix}" and "\bold x = (x, y, z)^T".

Search for general solution of this system using Gauss method (transforming matrix to row reduced form): "\begin{pmatrix}
1 & -2 & 3\\
-3 & 6 & 9
\end{pmatrix} \sim (II \rightarrow II + 3 I) \sim \begin{pmatrix}
1 & -2 & 3\\
0 & 0 & 0
\end{pmatrix}". As one can see, second row vanishes, so the rank of matrix "A" is 1.

Therefore, general solution to the system can be written as:"\bold x = (2 C_1 - 3 C_2; C_1; C_2)^T = C_1 (2; 1; 0)^T + C_2(-3; 0; 1)^T".

a) Basis of solution space consists of two vectors: "(2; 1; 0)^T" and "(-3; 0; 1)^T".

b) Dimension of the solution space is "2".